Find the domain and range of the following rational function.Use any notation.



f(x)=3/x-1 f(x)=2x/x-4 f(x)=x+3/5x-5 f(x)=2x+x/2x f(x)=(x2+4x+3)/x²-9

1
Expert's answer
2021-11-03T17:43:32-0400

"f(x)=\frac{3}{x-1}"

To find domain we set, Denominator=0

"x-1=0\\x=1"

When x=1 the function is undefined,


"\space f(x)=\frac{2x}{x-4}"

To find domain we set, Denominator=0

"\implies x-4=0\\\implies x=4"

When x=4 the function is undefined,

"\therefore" domain of "f(x)={x:x \ne 4}"

To find range,

"Y=\frac{2x}{x-4}\\"

"\implies \frac{1}{Y}=\frac{x-4}{2x}\\"

"\implies \frac{1}{Y}=\frac{1}{2}-\frac{2}{x}"

"\implies \frac{2}{x}=\frac{1}{2}-\frac{1}{Y}"

"\implies \frac{n}{2}=\frac{2Y}{Y-2}"

"\implies n=\frac{4Y}{Y-2}"

When, Y=2 the function is undefined

"\therefore" Range"=\{Y:Y\ne 2\}"


"f(x)=\frac{x+3}{5n-5}"

Denominator"=0\implies5n-5=0\\\implies n=1\\\therefore Domain=\{x|x\ne 1\}"

Now,

"Y=\frac{x+3}{5x-5}\\\implies5xY-5Y=x+3\\\implies x(5Y-1)=3+5Y\\n=\frac{3+5Y}{5Y-1}"

Set, "5Y-1=0\implies Y=\frac{1}{5}"

"\therefore" Range of "f(x)=\{Y|Y\ne \frac{1}{5}\}"


"f(x)=\frac{2x+2}{2x}"

To find domain we set, Denominator=0

"2x=0\\x=0"

When x=0 the function is undefined,


"f(x)=\frac{2x+4x+3}{2x^2-9}"

To find domain we set, Denominator=0

"x^2-9=0\\x^2=9\\x=3"

When x=3 the function is undefined,